# Causal Concept Embedding Models: Beyond Causal Opacity in Deep Learning

### v1

@@ C.3 Ablation study
Causal CEM $\text{Smiling (S)}&\leftarrow\epsilon_{0}\\ \text{Attractive (A)}&\leftarrow\text{Heavy\_Make}\\ \text{Mouth\_Slig (MS)}&\leftarrow\text{False}\\ \text{High\_Cheek (HC)}&\leftarrow\epsilon_{3}\\ \text{Wearing\_Li (WL)}&\leftarrow\text{Attractive}\\ \text{Heavy\_Make (HM)}&\leftarrow\text{Smiling}\wedge\text{High\_Cheek}\\ \text{Male}&\leftarrow\sim\text{Wearing\_Li}\wedge\sim\text{Heavy\_Make}\\ \text{Wavy\_Hair (WH)}&\leftarrow(\text{HC}\wedge\text{WL}\wedge\sim\text{Male})\vee(\text{HC}\wedge\text{WL}\wedge\sim\text{OF})\\ \text{Big\_Lips (BL)}&\leftarrow\text{Smiling}\wedge\text{High\_Cheek}\wedge\sim\text{Male}\wedge\sim\text{Oval\_Face}\\ \text{Oval\_Face (OF)}&\leftarrow\text{False}\\ \text{Makeup (M)}&\leftarrow\text{False}\\ \text{Fem\_Model}&\leftarrow(\text{M}\wedge\sim\text{S})\vee(\text{M}\wedge\sim\text{Male})\vee(\text{M}\wedge\text{HC}\wedge\sim\text{WH})\vee(\text{WL}\wedge\text{WH}\wedge\text{BL}\wedge\sim\text{S}\wedge\sim\text{HC})\\ \end{aligned}$
Causal CEM+CD $\text{Smiling}&\leftarrow\text{Mouth\_Slig}\\ \text{Attractive}&\leftarrow\text{Wearing\_Li}\\ \text{Mouth\_Slig}&\leftarrow\epsilon_{2}\\ \text{High\_Cheek}&\leftarrow\text{Smiling}\\ \text{Wearing\_Li}&\leftarrow\epsilon_{4}\\ \text{Heavy\_Make}&\leftarrow(\text{Attractive}\wedge\text{Wearing\_Li})\vee(\text{Wearing\_Li}\wedge\text{Oval\_Face})\\ \text{Male}&\leftarrow\sim\text{Wearing\_Li}\wedge\sim\text{Heavy\_Make}\\ \text{Wavy\_Hair}&\leftarrow\text{Makeup}\vee(\text{Attractive}\wedge\text{Wearing\_Li}\wedge\sim\text{Male})\\ \text{Big\_Lips}&\leftarrow\text{Makeup}\\ \text{Oval\_Face}&\leftarrow\text{Smiling}\wedge\text{Attractive}\wedge\text{Wearing\_Li}\wedge\sim\text{Big\_Lips}\\ \text{Makeup}&\leftarrow\text{False}\\ \text{Fem\_Model}&\leftarrow\text{Makeup}\\ \end{aligned}$
@@
Causal CEM 0.05 $88.24\pm 1.30$ $14.47\pm 2.86$ $00.00\pm 0$ 0.00 $44.37\pm 5.58$ $16.15\pm 11.25$ $00.00\pm 0$ 0.00
Causal CEM + CE 0.05 $89.43\pm 0.93$ $13.15\pm 2.31$ $00.00\pm 0$ 0.00 $39.53\pm 3.62$ $20.99\pm 8.19$ $00.00\pm 0$ 0.00
Causal CEM 0.2 $85.88\pm 3.31$ $11.64\pm 3.21$ $00.00\pm 0$ 0.00 $37.55\pm 8.77$ $16.32\pm 15.36$ $00.00\pm 0$ 0.00
Causal CEM + CE 0.2 $85.09\pm 2.78$ $16.39\pm 5.00$ $00.00\pm 0$ 0.00 $39.69\pm 10.97$ $31.15\pm 19.48$ $00.00\pm 0$ 0.00
Causal CEM 0 $87.86\pm 1.66$ $6.98\pm 2.95$ $00.00\pm 0$ 0.00 $18.16\pm 8.29$ $7.39\pm 4.10$ $00.00\pm 0$ 0.00
Causal CEM + CE 0 $87.04\pm 2.72$ $12.16\pm 2.92$ $00.00\pm 0$ 0.00 $33.15\pm 9.97$ $14.52\pm 14.29$ $00.00\pm 0$ 0.00
@@
Causal CEM + CE 0.05 $99.40\pm 0.15$ $12.85\pm 0.59$ $00.00\pm 0$ 0.00 $27.72\pm 3.66$ $28.95\pm 13.50$ $00.00\pm 0$ 0.00
Causal CEM 0.2 $98.80\pm 1.24$ $14.13\pm 3.39$ $00.00\pm 0$ 0.00 $37.59\pm 14.20$ $17.52\pm 13.41$ $00.00\pm 0$ 0.00
Causal CEM + CE 0.2 $99.30\pm 0.13$ $12.90\pm 0.51$ $00.00\pm 0$ 0.00 $34.01\pm 6.99$ $16.84\pm 6.79$ $00.00\pm 0$ 0.00
Causal CEM 0 $99.58\pm 0.12$ $6.89\pm 1.55$ $00.00\pm 0$ 0.00 $18.51\pm 5.14$ $12.11\pm 13.01$ $00.00\pm 0$ 0.00
Causal CEM + CE 0 $99.51\pm 0.05$ $5.67\pm 1.21$ $00.00\pm 0$ 0.00 $12.41\pm 1.82$ $15.15\pm 4.10$ $00.00\pm 0$ 0.00
@@
Causal CEM 0.05 $78.23\pm 0.45$ $2.17\pm 1.44$ $00.00\pm 0$ 0.00 $8.18\pm 4.38$ $0.04\pm 0.07$ $00.00\pm 0$ 0.00
Causal CEM + CE 0.05 $78.42\pm 0.42$ $5.48\pm 0.34$ $00.00\pm 0$ 0.00 $24.17\pm 1.32$ $1.24\pm 0.62$ $00.00\pm 0$ 0.00
Causal CEM 0.2 $77.49\pm 0.37$ $1.70\pm 0.98$ $00.00\pm 0$ 0.00 $8.95\pm 4.86$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Causal CEM + CE 0.2 $78.08\pm 0.39$ $6.15\pm 0.31$ $00.00\pm 0$ 0.00 $29.57\pm 1.03$ $0.82\pm 0.46$ $00.00\pm 0$ 0.00
Causal CEM 0 $77.42\pm 1.09$ $2.85\pm 0.44$ $00.00\pm 0$ 0.00 $13.06\pm 2.65$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Causal CEM + CE 0 $78.31\pm 0.36$ $4.64\pm 0.13$ $00.00\pm 0$ 0.00 $18.31\pm 2.61$ $1.25\pm 0.68$ $00.00\pm 0$ 0.00

### v4

@@ H.3 Ablation study
Causal CGM $\text{Smiling (S)}&\leftarrow\epsilon_{0}\\ \text{Attractive (A)}&\leftarrow\text{Heavy\_Make}\\ \text{Mouth\_Slig (MS)}&\leftarrow\text{False}\\ \text{High\_Cheek (HC)}&\leftarrow\epsilon_{3}\\ \text{Wearing\_Li (WL)}&\leftarrow\text{Attractive}\\ \text{Heavy\_Make (HM)}&\leftarrow\text{Smiling}\wedge\text{High\_Cheek}\\ \text{Male}&\leftarrow\sim\text{Wearing\_Li}\wedge\sim\text{Heavy\_Make}\\ \text{Wavy\_Hair (WH)}&\leftarrow(\text{HC}\wedge\text{WL}\wedge\sim\text{Male})\vee(\text{HC}\wedge\text{WL}\wedge\sim\text{OF})\\ \text{Big\_Lips (BL)}&\leftarrow\text{Smiling}\wedge\text{High\_Cheek}\wedge\sim\text{Male}\wedge\sim\text{Oval\_Face}\\ \text{Oval\_Face (OF)}&\leftarrow\text{False}\\ \text{Makeup (M)}&\leftarrow\text{False}\\ \text{Fem\_Model}&\leftarrow(\text{M}\wedge\sim\text{S})\vee(\text{M}\wedge\sim\text{Male})\vee(\text{M}\wedge\text{HC}\wedge\sim\text{WH})\vee(\text{WL}\wedge\text{WH}\wedge\text{BL}\wedge\sim\text{S}\wedge\sim\text{HC})\\ \end{aligned}$
Causal CGM+CD $\text{Smiling}&\leftarrow\text{Mouth\_Slig}\\ \text{Attractive}&\leftarrow\text{Wearing\_Li}\\ \text{Mouth\_Slig}&\leftarrow\epsilon_{2}\\ \text{High\_Cheek}&\leftarrow\text{Smiling}\\ \text{Wearing\_Li}&\leftarrow\epsilon_{4}\\ \text{Heavy\_Make}&\leftarrow(\text{Attractive}\wedge\text{Wearing\_Li})\vee(\text{Wearing\_Li}\wedge\text{Oval\_Face})\\ \text{Male}&\leftarrow\sim\text{Wearing\_Li}\wedge\sim\text{Heavy\_Make}\\ \text{Wavy\_Hair}&\leftarrow\text{Makeup}\vee(\text{Attractive}\wedge\text{Wearing\_Li}\wedge\sim\text{Male})\\ \text{Big\_Lips}&\leftarrow\text{Makeup}\\ \text{Oval\_Face}&\leftarrow\text{Smiling}\wedge\text{Attractive}\wedge\text{Wearing\_Li}\wedge\sim\text{Big\_Lips}\\ \text{Makeup}&\leftarrow\text{False}\\ \text{Fem\_Model}&\leftarrow\text{Makeup}\\ \end{aligned}$
@@
Causal CGM 0.05 $88.24\pm 1.30$ $14.47\pm 2.86$ $00.00\pm 0$ 0.00 $44.37\pm 5.58$ $16.15\pm 11.25$ $00.00\pm 0$ 0.00
Causal CGM + CE 0.05 $89.43\pm 0.93$ $13.15\pm 2.31$ $00.00\pm 0$ 0.00 $39.53\pm 3.62$ $20.99\pm 8.19$ $00.00\pm 0$ 0.00
Causal CGM 0.2 $85.88\pm 3.31$ $11.64\pm 3.21$ $00.00\pm 0$ 0.00 $37.55\pm 8.77$ $16.32\pm 15.36$ $00.00\pm 0$ 0.00
Causal CGM + CE 0.2 $85.09\pm 2.78$ $16.39\pm 5.00$ $00.00\pm 0$ 0.00 $39.69\pm 10.97$ $31.15\pm 19.48$ $00.00\pm 0$ 0.00
Causal CGM 0 $87.86\pm 1.66$ $6.98\pm 2.95$ $00.00\pm 0$ 0.00 $18.16\pm 8.29$ $7.39\pm 4.10$ $00.00\pm 0$ 0.00
Causal CGM + CE 0 $87.04\pm 2.72$ $12.16\pm 2.92$ $00.00\pm 0$ 0.00 $33.15\pm 9.97$ $14.52\pm 14.29$ $00.00\pm 0$ 0.00
@@
Causal CGM + CE 0.05 $99.40\pm 0.15$ $12.85\pm 0.59$ $00.00\pm 0$ 0.00 $27.72\pm 3.66$ $28.95\pm 13.50$ $00.00\pm 0$ 0.00
Causal CGM 0.2 $98.80\pm 1.24$ $14.13\pm 3.39$ $00.00\pm 0$ 0.00 $37.59\pm 14.20$ $17.52\pm 13.41$ $00.00\pm 0$ 0.00
Causal CGM + CE 0.2 $99.30\pm 0.13$ $12.90\pm 0.51$ $00.00\pm 0$ 0.00 $34.01\pm 6.99$ $16.84\pm 6.79$ $00.00\pm 0$ 0.00
Causal CGM 0 $99.58\pm 0.12$ $6.89\pm 1.55$ $00.00\pm 0$ 0.00 $18.51\pm 5.14$ $12.11\pm 13.01$ $00.00\pm 0$ 0.00
Causal CGM + CE 0 $99.51\pm 0.05$ $5.67\pm 1.21$ $00.00\pm 0$ 0.00 $12.41\pm 1.82$ $15.15\pm 4.10$ $00.00\pm 0$ 0.00
@@
Causal CGM 0.05 $78.23\pm 0.45$ $2.17\pm 1.44$ $00.00\pm 0$ 0.00 $8.18\pm 4.38$ $0.04\pm 0.07$ $00.00\pm 0$ 0.00
Causal CGM + CE 0.05 $78.42\pm 0.42$ $5.48\pm 0.34$ $00.00\pm 0$ 0.00 $24.17\pm 1.32$ $1.24\pm 0.62$ $00.00\pm 0$ 0.00
Causal CGM 0.2 $77.49\pm 0.37$ $1.70\pm 0.98$ $00.00\pm 0$ 0.00 $8.95\pm 4.86$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Causal CGM + CE 0.2 $78.08\pm 0.39$ $6.15\pm 0.31$ $00.00\pm 0$ 0.00 $29.57\pm 1.03$ $0.82\pm 0.46$ $00.00\pm 0$ 0.00
Causal CGM 0 $77.42\pm 1.09$ $2.85\pm 0.44$ $00.00\pm 0$ 0.00 $13.06\pm 2.65$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
Causal CGM + CE 0 $78.31\pm 0.36$ $4.64\pm 0.13$ $00.00\pm 0$ 0.00 $18.31\pm 2.61$ $1.25\pm 0.68$ $00.00\pm 0$ 0.00
H.4 Concept incompleteness and generalisation with a different number of concepts
Concept incompleteness does not have a strong impact on Causal CGM. As shown in the experiments with CLIP on CIFAR, we can automatically extract new concepts if the provided ones are insufficient. However, to quantify the impact of incompleteness on classification performances, we report an ablation study on a CBM benchmark (Marcinkevičs et al., 2024) specifically designed to assess concept incompleteness. This benchmark is a synthetic tabular dataset where we can control to what degree concepts explain the variance in the downstream task (see (Marcinkevičs et al., 2024) for details). In Figure 15 we observe that, as the concept incompleteness $i$ increases from $i=0$, where all relevant concepts are available, to $i=0.9$, where only $10\%$ of relevant concepts are available, the performance of CBMs drop, while that of black boxes, CEMs, and Causal CGMs do not thanks to the use of concept embeddings.
Figure 15: Label accuracy for different numbers of concepts with increasing concept incompleteness.
H.5 Scalability and computational complexity
Causal CGMs pays a small cost (at most quadratic in the number of class labels) to address causal transparency. In terms of the number of parameters, the endogenous predictor scales linearly with the number of concepts, as in standard CBMs. Regarding time complexity, the endogenous predictor scales quadratically (both in train and inference) with the number of concepts. This is because, during training, we model all concept dependencies, and, during testing, inference is performed on a DAG (a triangular adjacency matrix in the worst case). This is a small cost in exchange for interpretability (and not unique to our work). To better communicate our method’s scalability, we report the number of parameters and training and inference times as we increase the number of class labels on the same CBM benchmark used to assess incompleteness (Marcinkevičs et al., 2024). In Figure 16 we observe that the impact on the number of parameters is negligible as it is mostly due to the encoder. Here the encoder is a relatively small network, but it could be a billion-parameter in a Vision-lLanguage Model (VLM). Training and inference time are slightly higher than baselines, but if training includes a large encoder, say a VLM, the relative impact of the endogenous predictor would be significantly reduced (we did not have time to train a VLM from scratch).
Figure 16: Number of parameters, train and inference time.
H.6 $k$-unfolding ablation
In Causal CBM, $1$-step unfolding of the causal graph during training is equivalent to any k-step ($k>1$) unfolding as any step corresponds (theoretically and empirically) to repeating the same query on the same conditional probability table (CPT). Causal CGM learns the conditional probabilities of the variables given its parents (i.e. the CPT of the Bayesian network). Under full observability, conditional queries of the form $p(v|\text{ all but }v)$ are just inspections of the CPT and do not require any reasoning on the cyclical model (which alleviates us from setting a complex $k$-step semantics to unroll the cyclical model). The key observation is that the cyclical model served the purpose of templating multiple (any) acyclical graphs jointly without making assumptions on concept relations. During learning, the probabilities of the CPT (parameterized by NNs) are also subject to an acyclicity objective. After learning, the resulting graph is acyclic thus inference is performed via standard message passing along the directed acyclic graph. To show this experimentally, we report the results on our CBM benchmark using a different number of unfolding steps and considering different levels of concept completeness in Table 10. As expected, the results show that increasing the number of steps does not impact Causal CGM’s classification performances.
Incompleteness ratio
0.0 0.1 0.5 0.9
Unfolding K=1 $85.00\pm 0.14$ $83.70\pm 1.10$ $84.49\pm 0.57$ $85.51\pm 0.16$
Unfolding K=2 $83.78\pm 0.81$ $83.72\pm 0.85$ $85.26\pm 0.20$ $84.96\pm 0.14$
Unfolding K=3 $85.00\pm 0.34$ $83.10\pm 0.98$ $84.30\pm 0.47$ $85.32\pm 0.05$
Table 10: Causal CGM’s accuracy with $k=\{1,2,3\}$-unfolding and higher concept incompleteness.
Appendix I Answering counterfactual queries with Causal Concept Graph Models
In our experiments, we use Causal CGMs to compute Pearl’s Probability of Sufficiency and Necessity (PNS), which is a prototypical example of a counterfactual query (Tian & Pearl, 2000). In general, the computation of counterfactual queries requires three key ingredients  (Pearl, 2009): (i) a causal graph, (ii) structural rules predicting an endogenous variable given its parents’ values, and (iii) information about the exogenous variables distribution, which for us is encoded in the embeddings. In causal CGM, all these three ingredients are available, allowing the computation of both identifiable and unidentifiable causal queries. For the latter, as is the case of PNS, the result is given in terms of an interval exactly as it happens in causal inference with structural causal models (with probabilistic quantification over exogenous variables). Figure 17 illustrates the upper boundaries of the PNS tables obtained using Causal CGM across the four datasets. Additionally, Table 11 provides an example of counterfactuals generated on the CelebA dataset, where we first intervene on the lipstick concept, followed by an intervention on the makeup concept.
Figure 17: The upper boundary of the PNS computed on the Causal CGM models on the four different datasets. The PNS is computed between each couple of nodes that are directly or indirectly connected in the graph.
Table 11: Examples of counterfactuals generated on CelebA obtained via do-interventions (intervened variables are marked in red, changed variables are underlined).
Endogenous variables’ activations
Initial Predicted State Smiling, Not Attractive, Mouth Slig, High Cheek, Not Wearing Li, Not Heavy Make,
Not Wavy Hair, Not Big Lips, Not Oval Face, Not Makeup, Not Fem Model
What if I would wear lipstick? Smiling, Attractive, Mouth Slig, High Cheek, Wearing Li, Heavy Make,
Not Wavy Hair, Not Big Lips, Not Oval Face, Not Makeup, Not Fem Model
What if I would wear lipstick and also makeup? Smiling, Attractive, Mouth Slig, High Cheek, Wearing Li, Heavy Make,
Not Wavy Hair, Big Lips, Not Oval Face, Makeup, Fem Model
Appendix J Causal Concept Graph Models address “causal opacity”, not “causal discovery”
Our work is based on a fundamental yet subtle distinction between causal opacity/transparency and causal discovery that is poorly understood in the ML literature. Although related, the two problems are different in nature: Causal opacity/transparency refers to the problem of understanding and manipulating the causal structure of a model’s decision-making process, and it thus is a specific instance of what is usually referred to as the opacity or black-box problem. Causal discovery, on the other hand, is not concerned with how the model is but rather with how the world (Boge, 2022), i.e., it concerns determining the causal structure of the generating mechanism beyond data. This distinction applies also to the notion of causal understanding. Indeed, we can distinguish between two forms of causal understanding depending on whether the object of one’s understanding is the model’s decision-making process or the world/data-generating mechanism. Notice that causal opacity (and the related form of causal understanding) neither implies nor requires causal discovery. The causal structure of the model’s decision-making process does not necessarily mirror the causal structure of the world. Of course, one could argue that a good model should always resemble the world’s causal structure, but this is not always true or even wanted (e.g., if the world is biased, we may want to construct an unbiased model that deviates from the data-generating process of its training data).
Appendix K Relations with causal representation learning
If we assume a strict definition of causal representation learning (CLR) as any method that aims at learning high-level representations of the causal-generating mechanisms of data (Schölkopf et al., 2021), then Causal CGMs are not an example of CRL. Causal CGMs aims not to discover causal mechanisms beyond data but to build models whose decision-making processes follow a causal structure that users can visualize, interpret, and manipulate. This requires the model to use high-level variables (concepts) that are meaningful for a user (as opposed to, say, an image’s pixels) and to clearly display the causal relationships between these variables (e.g., via a casual graph).
Regardless, there are similarities between Causal CGMs and CRL. Both rely on Pearl’s formalism of SCMs and use high-level interpretable variables and causal graphs to model causal dependencies. In the case of Causal CGMs, however, the learned causal structure may not necessarily mirror the causal structure of the data-generating mechanism, as it is required for CRL models.
Appendix L Comparing different decision making processes with Causal Concept Graph Models
Causal CGMs are not identifiable in the sense that we can obtain different causal graphs by training the model over the same dataset. This is similar to obtaining different parameter values and embeddings when training standard deep-learning black boxes on the same data. Similarly, the training of causal CGMs can result in different causal decision-making structures for the same task. The difference is that in the case of the black box, we cannot compare the different models because they are (causally) opaque. In the case of Causal CGMs, we can inspect the multiple models obtained and verify them against desirable properties (e.g., fairness), hence adopting the best one.